A mass of 2.3 kg is initially moving at 5 m/s. A constant force of 15N is applied to the mass for 6 seconds. Find (a) the initial momentum, (b) final momentum.


1
Expert's answer
2021-12-22T14:09:00-0500

(a)

"p_i=mv_i=2.3\ kg\times5\ \dfrac{m}{s}=11.5\ \dfrac{kg\times m}{s}."

(b)

"F\Delta t=\Delta p=p_f-p_i,""p_f=F\Delta t+p_i,""p_f=15\ N\times6\ s+11.5\ \dfrac{kg\times m}{s}=101.5\ \dfrac{kg\times m}{s}"